OpenStudy (anonymous):

Two trains leave stations 324miles apart at the same time and travel toward each other. One train travels at 85 miles per hour while the other travels at 95 miles per hour. How long will it take for the two trains to meet?

OpenStudy (anonymous):

relative velocity of both trains=85+95 (towards each other) so total distance=324 and hence relative time that both trains would take to meet= distance / relative speed =324/180 =1.8hr